OpenStudy (anonymous):

The length of a rectangle is 3 more than twice the width. The area of the rectangle is 119 square inches. What are the dimensions of the rectangle? If x = the width of the rectangle, which of the following equations is used in the process of solving this problem? 2x^2 + 3x - 119 = 0 3x^2 + 3x - 119 = 0 6x^2 - 119 = 0

OpenStudy (anonymous):

\[Area=l*w\]

OpenStudy (anonymous):

so it is the 1st one

OpenStudy (anonymous):

\[(2x+3)x=119\]\[2x^2+3x=119\] \[2x^2+3x-119=0\] yup thats right.

OpenStudy (anonymous):

ok thanks